# spacing=nonfrench

### v2

@@ 6.2 Limitations and Future Work
For the limitations, as discussed in Section 3, the per-query $\epsilon$-LDP privacy guarantee only holds under the assumption that each data point in ${\mathcal{D}_{\rm priv}}$ is i.i.d.. If they are not i.i.d., our privacy guarantee will be weakened by their mutual information. Many real-world applications may use data de-duplication technologies to maintain the i.i.d. assumption for data management, ensuring data integrity and preventing redundancy. For instance, cloud storage services such as Amazon S3 identify and delete duplicated objects to optimize user’s storage space [67]. Entities and businesses such as credit bureaus and e-commerce platforms use de-duplication techniques to merge and resolve records based on attributes (e.g., names, birth dates, and addresses), even when minor differences exist.
@@
As discussed in Section 4.4, LDPKiT is compatible with advanced training strategies, such as active learning, that expedite model training and help training set size reduction. Active learning strategies can also be applied to prune queries in ${\mathcal{D}_{\rm infer}}$ that are less significant. We plan to study the effect of such strategies on LDPKiT in the future. Furthermore, since ${\mathcal{M}_{\rm L}}$ is trained on noisy data, LDPKiT is inherently immune to membership inference attacks if ${\mathcal{M}_{\rm L}}$ is ever leaked. We leave the analyses of membership inference and attribute inference attack success rates as future work.
Due to time constraints, we only tested on supervised learning, specifically classification tasks. In the future, we may extend the evaluation to regression tasks or unsupervised clustering tasks. Furthermore, while the empirical analyses in Section 4 focus on image benchmarks, extending LDPKiT to other modalities, such as text, is part of our future work.
@@
Also, our privacy-preserving queries in this paper refer to LDPKiT-generated noised queries with LDPKiT-Rand or LDPKiT-Sup utilizing $\epsilon$-LDP privacy mechanism. We can generate such queries in other ways. For instance, LDPKiT is compatible with other noise mechanisms such as ($\epsilon$, $\delta$)-LDP with the Gaussian mechanism. In addition, we can also compose our privacy-preserving ${\mathcal{D}_{\rm priv}}$ by generating synthetic queries using a Generative Adversarial Network [68].
7 Conclusion
LDPKiT is an inference framework designed to protect the privacy of sensitive data when using privacy-sensitive yet potentially malicious cloud services. LDPKiT achieves privacy protection by introducing LDP noise into the data before transmission to the cloud model for inference, ensuring privacy even if the platform or model is compromised. Moreover, LDPKiT mitigates the accuracy loss with a two-layer noise injection mechanism, e.g., LDPKiT-Rand and LDPKiT-Sup, leveraging the post-processing property of LDP. The base noise ensures privacy protection, while the second layer of noise helps utility recovery by generating sufficient data points that closely approximate the target distribution of the private data for local training. The key insight is that effective yet privacy-preserving knowledge transfer requires the user to query a sufficient amount of privacy-protected noised data points that have a closer distribution to the actual private data. Superimposition in LDPKiT-Sup creates such a dataset that is more representative than LDPKiT-Rand which adds unrelated noise. The experimental results show that LDPKiT-Sup successfully recovers prediction accuracy on private data during training while preserving privacy. Furthermore, quantitative analysis such as latent space analysis confirms the accuracy improvements. LDPKiT-Sup has greater benefits with higher noise levels, corresponding to stronger privacy guarantees. LDPKiT employs knowledge transfer and model extraction techniques while adhering to cloud service’s terms of use by limiting use to non-commercial purposes and ensuring derived models do not compete commercially. Additionally, quantitative evaluations confirm that the knowledge transfer in LDPKiT does not facilitate adversarial model extraction attacks, reinforcing its ethical and secure design.
8 Ethics Consideration
The author(s) attest that they read and follow the USENIX Security ’25 Ethics Guidelines. This research is conducted with a strong commitment to ethical integrity. This discussion of ethics consideration focuses on identifying stakeholders, assessing potential risks, implementing mitigations, and justifying key decisions. The primary stakeholders include service providers offering the cloud service platform, end users whose private data may be transmitted to the cloud and face potential privacy risks, and research team member(s). Given the use of knowledge transfer and model extraction techniques in this research, the experiments could potentially conflict with the Terms of Service set by the service providers. However, the authors affirm that the extraction techniques employed do not violate these Terms of Service and do not constitute an adversarial model extraction attack (i.e., model theft). Further details are provided in Section 6.1 of the paper.
Additionally, the authors confirm that the research does not compromise end users’ private data. Instead, it focuses on developing a privacy-preserving framework that safeguards user data privacy in the context of cloud ML applications.
Lastly, the research activities have been conducted with care to ensure there is no negative impact on the wellbeing of research team member(s).
9 Open science
The author(s) attest that they comply with the USENIX Security ’25 Open Science Policies. The author(s) will make the research artifacts available to the public upon paper acceptance. Specifically, the source code and scripts will be shared to ensure the availability, functionality and reproducibility of the research artifacts, facilitating further research and development in this area.
@@ 9 Open science
Appendix A Proof of $\epsilon$-LDP with the Laplacian mechanism and Post-Processing Property
@@
We discuss the proof aforementioned in Section 3 here: We prove that the base noise injection scheme (i.e., the first layer of noise) that generates ${\mathcal{D}_{\rm protected}}$ satisfies the definition of ${\epsilon}$-LDP with the Laplacian mechanism in Theorem A.1.
@@
Let $v$ be the original data and $f(v)$ be the query and computation function performed on the data with function sensitivity, $\Delta_{f}=\max_{v_{1},v_{2}}\|f(v_{2})-f(v_{1})\|_{1}$. We define the randomization algorithm $\mathcal{A}$ with Laplacian mechanism such that for any input value $v$, $\mathcal{A}=f(v)+\mathcal{Z}$, where $\mathcal{Z}$ is sampled from the Laplacian distribution $L(z,\lambda)$ with scaling factor $\lambda$ set to $\frac{\Delta_{f}}{\epsilon}$.
@@
By substituting the PDF of Laplacian distribution’s new expression into the Equation 1, we get
@@
Then, we prove that our noise generation algorithms, LDPKiT-Rand and LDPKiT-Sup, satisfy the post-processing property of LDP and hence provide the same level of $\epsilon$-LDP privacy guarantee on ${\mathcal{D}_{\rm infer}}$ as ${\mathcal{D}_{\rm protected}}$.
@@
Our noise injection mechanism LDPKiT-Rand satisfies ${\epsilon}$-LDP.
@@
For each data point $d_{i}\in\mathcal{D}_{\text{priv}}$, where $i=1,2,\dots,|{\mathcal{D}_{\rm priv}}|$, apply the Laplacian mechanism (See Definition 2.2) to obtain the first noise-added data point:
@@
By Definition 2.1, this initial mechanism satisfies $\epsilon$-LDP.
As for the post-processing step in LDPKiT-Rand, for each $\tilde{d}_{i}^{(1)}$, we generate $|{\mathcal{D}_{\rm priv}}|-1$ distinct random versions by adding an additional layer of Laplacian noise with the same scale:
$\tilde{d}_{i,j}^{(2)}=\tilde{d}_{i}^{(1)}+\eta_{i,j}^{(2)},\quad\text{for }j=1,2,\dots,|{\mathcal{D}_{\rm priv}}|-1,$
where $\eta_{i,j}^{(2)}$ are independent random variables drawn from the Laplacian distribution with scaling factor $\lambda=\frac{\Delta_{f}}{\epsilon}$.
@@
Hence, the protected inference dataset ${\mathcal{D}_{\rm infer}}$ at most consists of all $|{\mathcal{D}_{\rm priv}}|\cdot(|{\mathcal{D}_{\rm priv}}|-1)$ data points $\tilde{d}_{i,j}^{(2)}$:
${\mathcal{D}_{\rm infer}}=\left\{\tilde{d}_{i,j}^{(2)}\;\middle|\;\begin{array}[]{l}i=1,2,\dots,|{\mathcal{D}_{\rm priv}}|;\\ j=1,2,\dots,|{\mathcal{D}_{\rm priv}}|-1\end{array}\right\}.$
By Definition 2.3, since the initial mechanism $\mathcal{A}$ satisfies $\epsilon$-LDP, any function $g$ that processes its output without accessing the original data preserves the $\epsilon$-LDP guarantee. The addition of Laplacian noise to $\tilde{d}_{i}^{(1)}$ is a randomized function $g$ that depends only on $\tilde{d}_{i}^{(1)}$ and independent random noise $\eta_{i,j}^{(2)}$. Hence, the composite mechanism $g(\mathcal{A}(d_{i}))$ satisfies:
$\Pr\left[g(\mathcal{A}(v_{1}))\in\mathcal{T}\right]\leq e^{\epsilon}\Pr\left[g(\mathcal{A}(v_{2}))\in\mathcal{T}\right],$ (5)
for all values $v_{1},v_{2}$ and all subsets $\mathcal{T}\subseteq\text{Range}(g(\mathcal{A}))$.
@@
Our noise injection mechanism LDPKiT-Sup satisfies ${\epsilon}$-LDP.
@@
The base noise addition mechanism $\mathcal{A}$ applies the Laplacian mechanism (See Definition 2.2) to obtain the noise-added data point:
@@
By Definition 2.1, the base noise addition mechanism that outputs $\tilde{d}_{i}$ satisfies $\epsilon$-LDP.
We define a post-processing function $g$ that takes the set of noise-added data points $\{\tilde{d}_{1},\tilde{d}_{2},\dots,\tilde{d}_{|}{\mathcal{D}_{\rm priv}}|\}$ and outputs all possible ordered pairs:
[…]
The total number of data points in ${\mathcal{D}_{\rm infer}}$ is $|{\mathcal{D}_{\rm priv}}|\cdot(|{\mathcal{D}_{\rm priv}}|-1)$.
By Definition 2.3, if the base noise addition mechanism $\mathcal{A}$ satisfies $\epsilon$-LDP, then any function $g$ applied to its output preserves the $\epsilon$-LDP guarantee.
@@
$\Pr\left[g(\mathcal{A}(v_{1}))\in\mathcal{T}\right]\leq e^{\epsilon}\Pr\left[g(\mathcal{A}(v_{2}))\in\mathcal{T}\right],$ (6)
for all values $v_{1},v_{2}$ and all subsets $\mathcal{T}\subseteq\text{Range}(g(\mathcal{A}))$.
@@
(a) PathMNIST
(b) Fashion-MNIST
(c) SVHN
Figure 14: Impact of $\epsilon$ values on MobileNetV2 (${\mathcal{M}_{\rm L}}$)’s accuracies on ${\mathcal{D}_{\rm priv}}$ versus ${\mathcal{D}_{\rm val}}$ with LDPKiT-Sup.
(a) PathMNIST
(b) Fashion-MNIST
(c) SVHN
Figure 15: Impact of $\epsilon$ values on MobileNetV2 (${\mathcal{M}_{\rm L}}$)’s accuracies on ${\mathcal{D}_{\rm priv}}$ versus ${\mathcal{D}_{\rm val}}$ with LDPKiT-Rand.
(a) Fashion-MNIST ($\epsilon=1.5$)
(b) SVHN ($\epsilon=2.0$)
(c) SVHN ($\epsilon=1.25$)
(d) PathMNIST ($\epsilon=7.0$)
Table 5: Final accuracies on ${\mathcal{D}_{\rm priv}}$ and ${\mathcal{D}_{\rm val}}$ of SVHN.
Model $\epsilon$ Value Accuracy on ${\mathcal{D}_{\rm priv}}$ (%) Accuracy on ${\mathcal{D}_{\rm val}}$ (%)
@@
SIDP is inference accuracy of ResNet-152 (${\mathcal{M}_{\rm R}}$) on ${\mathcal{D}_{\rm protected}}$ without utility trade-off mitigation. The values recorded in parentheses are the standard deviations of the accuracies.
Table 6: Final accuracies on ${\mathcal{D}_{\rm priv}}$ and ${\mathcal{D}_{\rm val}}$ of Fashion-MNIST.
Model $\epsilon$ Value Accuracy on ${\mathcal{D}_{\rm priv}}$ (%) Accuracy on ${\mathcal{D}_{\rm val}}$ (%)
@@
SIDP is inference accuracy of ResNet-152 (${\mathcal{M}_{\rm R}}$) on ${\mathcal{D}_{\rm protected}}$ without utility trade-off mitigation. The values recorded in parentheses are the standard deviations of the accuracies.
Table 7: Final accuracies on ${\mathcal{D}_{\rm priv}}$ and ${\mathcal{D}_{\rm val}}$ of PathMNIST.
Model $\epsilon$ Value Accuracy on ${\mathcal{D}_{\rm priv}}$ (%) Accuracy on ${\mathcal{D}_{\rm val}}$ (%)
@@
SIDP is inference accuracy of ResNet-152 (${\mathcal{M}_{\rm R}}$) on ${\mathcal{D}_{\rm protected}}$ without utility trade-off mitigation. The values recorded in parentheses are the standard deviations of the accuracies.
Table 8: Sensitivity Analysis on $|{\mathcal{D}_{\rm priv}}|$ and $|{\mathcal{D}_{\rm infer}}|$ with ResNet-18 (${\mathcal{M}_{\rm L}}$) (Continued from Table 3).

### v3

@@ VI-B Limitations and Future Work
As discussed in Section III, our per-query $\epsilon$-LDP privacy guarantee relies on the assumption that each data point in $\mathcal{D}_{\rm priv}$} is i.i.d.. If the data points are not i.i.d., the privacy guarantees can be weakened due to their mutual information. In practice, many real-world applications leverage data de-duplication techniques to maintain or approximate the i.i.d. assumption. For instance, cloud storage services such as Amazon S3 implement de-duplication to identify and remove duplicate files, optimizing user storage space [90]. Similarly, entities such as credit bureaus and e-commerce platforms employ de-duplication methods to consolidate and accurately resolve records based on shared attributes, including names, birth dates, addresses, and transaction histories, even when slight variations exist. Additionally, in the domain of generative AI, training LLMs often involves extensive data de-duplication to improve computational efficiency, enhance model performance, and reduce the risk of privacy leakage by removing redundant or duplicated content from training datasets [91, 92, 93, 94].
In our experiments, we assume that the ML service model is honest-but-curious; if the model does not always provide correct answers to the inference queries, an increased number of queries and higher cost would be incurred.
Currently, LDPKiT only supports the image modality, where superimposition intuitively translates to pixel-level operations. Although textual data does not have a counterpart for superimposition, several conceptual approaches can be explored. These include embedding-level averaging, in which two textual embeddings are blended; token-level mixing, which involves the interleaving of tokens from two samples; and semantic-level merging, such as combining two sentences into a coherent paraphrased summary. Each interpretation may require a distinct adaptation of the LDP scheme. We leave the exploration of these textual superimposition methods as a direction for future research.
@@
As discussed in Section IV-D, LDPKiT is compatible with advanced training strategies, such as active learning, which expedite model training and reduce the required training set size. Active learning strategies can also be applied to prune queries in $\mathcal{D}_{\rm infer}$} that are less significant. We plan to study the effect of such strategies on LDPKiT in the future. Furthermore, since $\mathcal{M}_{\rm L}$} is trained on noisy data, LDPKiT is inherently immune to membership inference attacks if $\mathcal{M}_{\rm L}$} is ever leaked. We leave the analyses of membership inference and attribute inference attack success rates as future work.
@@
Further, our privacy-preserving queries in this paper refer to LDPKiT-generated noised queries with LDPKiT-Rand or LDPKiT-Sup utilizing $\epsilon$-LDP privacy mechanism. We can generate such queries in other ways. For instance, LDPKiT is compatible with other noise mechanisms such as ($\epsilon$, $\delta$)-LDP with the Gaussian mechanism [32]. In addition, we can also compose our privacy-preserving $\mathcal{D}_{\rm priv}$} by generating synthetic queries using a generative model such as a diffusion model or an LLM. We leave a detailed analysis of these alternate techniques for future work.
We also extended the evaluation of LDPKiT to the ImageNet1K dataset [95], where all images were resized to 256×256 pixels and subsequently center-cropped to 224×224 pixels, consistent with the standard preprocessing procedure in [96]. For reference, the official torchvision ResNet-18 model reports a top-1 accuracy of 69.758% on ImageNet1K [96]. In our evaluation, a randomly selected subset of 15 classes was used. Under this setting, LDPKiT-Sup with ResNet-18 ($\mathcal{M}_{\rm L}$}) achieved a top-1 accuracy of 63% with 15,000 $\mathcal{D}_{\rm priv}$} samples at $\epsilon=5$, whereas the baseline SIDP attained only 19% accuracy under identical conditions. Further enhancing the performance of LDPKiT on larger datasets is left for future work.
VII Conclusion
LDPKiT provides a novel privacy-preserving inference framework designed to protect sensitive user data when interacting with powerful yet potentially malicious ML model providers. By incorporating LDP noise into inference queries, LDPKiT protects privacy even if service platforms or models are compromised. The key contribution of LDPKiT is the integration of LDP with a two-layer data augmentation mechanism (LDPKiT-Rand and LDPKiT-Sup), leveraging LDP’s post-processing property to enhance utility without compromising privacy. The first layer provides the baseline LDP privacy protection, and the second layer recovers utility with a larger augmented dataset generated for effective local model extraction. Specifically, LDPKiT-Sup employs a novel superimposition-based augmentation technique to produce more representative data points that better approximate the original private data distribution compared to the random-noise strategy in LDPKiT-Rand. The extensive evaluation results show that LDPKiT-Sup successfully recovers prediction accuracy on private data while preserving privacy. It has greater benefits at higher noise levels, corresponding to stronger privacy guarantees. Furthermore, LDPKiT explicitly acknowledges the economic and ethical considerations: it preserves service providers’ revenue models with user-issued queries and ensures that the extracted models are non-stealthy and non-competitive, complying with providers’ terms of use. In doing so, LDPKiT maintains an ethical and secure design that upholds the principles of responsible data usage and privacy by design.
VIII LLM Usage Considerations
The authors used generative AI-based tools, including OpenAI’s ChatGPT [97] and Grammarly [98], to correct typos, grammatical errors, and awkward phrasings in the preparation of this manuscript. All AI-generated outputs were inspected by the authors to ensure accuracy and originality.
Appendix A Appendix
(a) Label 2
(b) Label 1
Figure 8: Samples of noised SVHN data with $\epsilon$ set to 1.5 (left) and 1.25 (right).
(a) Label 0 (T-shirt/top)
(b) Label 6 (shirt)
Figure 9: Samples of noised Fashion-MNIST data with $\epsilon$ set to 2.0 (left) and 1.5 (right).
(a) Label 0 (adipose)
(b) Label 0 (adipose)
Figure 10: Samples of noised PathMNIST data with $\epsilon$ set to 10.0 (left) and 7.0 (right).
@@ Appendix A Appendix
A-A Proof of $\epsilon$-LDP with the Laplacian mechanism and Post-Processing Property
@@
We discuss the proof aforementioned in Section III here: We prove that the base noise injection scheme (i.e., the first layer of noise) generating $\mathcal{D}_{\rm protected}$} satisfies the definition of $\epsilon$}-LDP with the Laplacian mechanism in Theorem A.1.
@@
Let $v$ be the original data and $f(v)$ be the query and computation function performed on the data with function sensitivity, $\Delta_{f}=\max_{v_{1},v_{2}}\|f(v_{2})-f(v_{1})\|_{1}$. We define the randomization algorithm $\mathcal{A}$ with Laplacian mechanism such that for any input value $v$, $\mathcal{A}=f(v)+\mathcal{Z}$, where $\mathcal{Z}$ is sampled from the Laplacian distribution $\mathcal{L}(z,\lambda)$ with scaling factor $\lambda$ set to $\frac{\Delta_{f}}{\epsilon}$.
@@
By substituting the PDF of the Laplacian distribution’s new expression into Equation 1, we get
@@
Then, we prove that our data augmentation algorithms, LDPKiT-Rand and LDPKiT-Sup, satisfy the post-processing property of LDP and hence provide the same level of $\epsilon$-LDP privacy guarantee on $\mathcal{D}_{\rm infer}$} as $\mathcal{D}_{\rm protected}$}. For convenience, we assume the dataset used for querying, $\mathcal{D}_{\rm infer}$}, is the maximum possible augmented inference dataset generated from $\mathcal{D}_{\rm priv}$} without post-selection per definitions in Section III-B.
@@
Our data augmentation mechanism LDPKiT-Rand satisfies $\epsilon$}-LDP.
@@
For each data point $d_{i}\in\mathcal{D}_{\text{priv}}$, where $i=1,2,\dots,|{\mathcal{D}_{\rm priv}}|$ (note that $|{\mathcal{D}_{\rm protected}}|=|{\mathcal{D}_{\rm priv}}|$), apply the Laplacian mechanism (See Definition II.2) to obtain the first noise-added data point:
@@
By Definition II.1, this initial mechanism satisfies $\epsilon$-LDP.
As for the post-processing step in LDPKiT-Rand, for each $\tilde{d}_{i}^{(1)}$, we generate $|{\mathcal{D}_{\rm protected}}|-1$ distinct random versions by adding an additional layer of Laplacian noise with the same scale:
$\tilde{d}_{i,j}^{(2)}=\tilde{d}_{i}^{(1)}+\eta_{j}^{(2)},$
where $i=1,2,\dots,|{\mathcal{D}_{\rm protected}}|,j=1,2,\dots,|{\mathcal{D}_{\rm protected}}|-1$ and $\eta_{j}^{(2)}$ are independent random variables drawn from the Laplacian distribution with scaling factor $\lambda=\frac{\Delta_{f}}{\epsilon}$.
@@
Hence, the protected inference dataset $\mathcal{D}_{\rm infer}$} at most consists of all $|{\mathcal{D}_{\rm protected}}|\cdot(|{\mathcal{D}_{\rm protected}}|-1)$ data points $\tilde{d}_{i,j}^{(2)}$:
${\mathcal{D}_{\rm infer}}=\left\{\tilde{d}_{i,j}^{(2)}\;\middle|\;\begin{array}[]{l}i=1,2,\dots,|{\mathcal{D}_{\rm protected}}|;\\ j=1,2,\dots,|{\mathcal{D}_{\rm protected}}|-1\end{array}\right\}.$
By Definition II.3, since the initial mechanism $\mathcal{A}$ satisfies $\epsilon$-LDP, any function $g$ that processes its output without accessing the original data preserves the $\epsilon$-LDP guarantee. The addition of Laplacian noise to $\tilde{d}_{i}^{(1)}$ is a randomized function $g$ that depends only on $\tilde{d}_{i}^{(1)}$ and independent random noise $\eta_{j}^{(2)}$. Hence, the composite mechanism $g(\mathcal{A}(d_{i}))$ satisfies:
$\Pr\left[g(\mathcal{A}(d_{1}))\in\mathcal{T}\right]\leq e^{\epsilon}\Pr\left[g(\mathcal{A}(d_{2}))\in\mathcal{T}\right],$ (5)
for all values $d_{1},d_{2}$ and all subsets $\mathcal{T}\subseteq\text{Range}(g(\mathcal{A}))$.
@@
Our data augmentation mechanism LDPKiT-Sup satisfies $\epsilon$}-LDP.
@@
The base noise addition mechanism $\mathcal{A}$ applies the Laplacian mechanism (See Definition II.2) to obtain the noise-added data point:
@@
By Definition II.1, the base noise addition mechanism that outputs $\tilde{d}_{i}$ satisfies $\epsilon$-LDP.
We define a post-processing function $g$ that takes the set of noise-added data points $\{\tilde{d}_{1},\tilde{d}_{2},\dots,\tilde{d}_{|}{\mathcal{D}_{\rm protected}}_{|}\}$ and outputs the average of all possible ordered pairs:
${\mathcal{D}_{\rm infer}}=g\left(\{\tilde{d}_{i}\}\right)=\left\{\begin{aligned} &avg(\tilde{d}_{i},\tilde{d}_{j})\;\mid\;i\neq j,\\ &1\leq i,j\leq|{\mathcal{D}_{\rm protected}}|\end{aligned}\right\}.$
The size of $\mathcal{D}_{\rm infer}$} is $|{\mathcal{D}_{\rm protected}}|\cdot(|{\mathcal{D}_{\rm protected}}|-1)$.
By Definition II.3, if the base noise addition mechanism $\mathcal{A}$ satisfies $\epsilon$-LDP, then any function $g$ applied to its output preserves the $\epsilon$-LDP guarantee.
@@
$\Pr\left[g(\mathcal{A}(d_{1}))\in\mathcal{T}\right]\leq e^{\epsilon}\Pr\left[g(\mathcal{A}(d_{2}))\in\mathcal{T}\right],$ (6)
for all values $d_{1},d_{2}$ and all subsets $\mathcal{T}\subseteq\text{Range}(g(\mathcal{A}))$.
@@
TABLE V: Final accuracies on $\mathcal{D}_{\rm priv}$} and $\mathcal{D}_{\rm val}$} of SVHN.
Model $\epsilon$
Accuracy on $\mathcal{D}_{\rm priv}$} (%)
Accuracy on $\mathcal{D}_{\rm val}$} (%)
@@
SIDP is the inference accuracy of ResNet-152 ($\mathcal{M}_{\rm R}$}) on $\mathcal{D}_{\rm protected}$} without utility trade-off mitigation. The values recorded in parentheses are the standard deviations of the accuracies.
TABLE VI: Final accuracies on $\mathcal{D}_{\rm priv}$} and $\mathcal{D}_{\rm val}$} of Fashion-MNIST.
Model $\epsilon$
Accuracy on $\mathcal{D}_{\rm priv}$} (%)
Accuracy on $\mathcal{D}_{\rm val}$} (%)
@@
SIDP is the inference accuracy of ResNet-152 ($\mathcal{M}_{\rm R}$}) on $\mathcal{D}_{\rm protected}$} without utility trade-off mitigation. The values recorded in parentheses are the standard deviations of the accuracies.
TABLE VII: Final accuracies on $\mathcal{D}_{\rm priv}$} and $\mathcal{D}_{\rm val}$} of PathMNIST.
Model $\epsilon$
Accuracy on $\mathcal{D}_{\rm priv}$} (%)
Accuracy on $\mathcal{D}_{\rm val}$} (%)
@@
SIDP is the inference accuracy of ResNet-152 ($\mathcal{M}_{\rm R}$}) on $\mathcal{D}_{\rm protected}$} without utility trade-off mitigation. The values recorded in parentheses are the standard deviations of the accuracies.
A-B Examples of $\mathcal{D}_{\rm infer}$} data with LDP noise
Figures 8, 9, and 10 present comparisons between the original and noised data samples under different levels of $\epsilon$-LDP noise. They also illustrate the impact of our two privacy-preserving data augmentation mechanisms, LDPKiT-Rand and LDPKiT-Sup.
(a) Fashion-MNIST ($\epsilon=1.5$)
(b) SVHN ($\epsilon=2.0$)
(c) SVHN ($\epsilon=1.25$)
(d) PathMNIST ($\epsilon=7.0$)
Figure 12: Qualitative results of image reconstruction with diffusion models. (Top row: privacy-preserving data generated with LDPKiT-Sup. Bottom row: applying the reconstruction attack on data protected by LDPKiT-Sup).
Figure 13: Sample reconstructed images demonstrating the effectiveness of the denoiser on SVHN and PathMNIST when little noise is added (e.g., $\epsilon$ = 15.0 and 20.0).
A-C Hyperparameter choices and dataset preparation
In this section, we document hyperparameter choices and dataset splits for the experiments in Section IV. We use three image benchmarks for evaluation: SVHN with 10 classes of 32x32x3 cropped street view house number images, Fashion-MNIST with 10 classes of 28x28 greyscale fashionable clothing images, and PathMNIST with 9 classes of 28x28x3 medical images of pathology. We train $\mathcal{M}_{\rm R}$} with a learning rate of 0.1 for 200 epochs on Fashion-MNIST and SVHN and 10 epochs on PathMNIST. $\mathcal{M}_{\rm R}$} is trained on 35k data points for Fashion-MNIST, 48,257 data points for SVHN and 89,996 data points for PathMNIST. $\mathcal{D}_{\rm priv}$} and $\mathcal{D}_{\rm val}$} are split from the remaining data points unseen by $\mathcal{M}_{\rm R}$}, where $\mathcal{D}_{\rm priv}$} is used to train and evaluate $\mathcal{M}_{\rm L}$}, and $\mathcal{D}_{\rm val}$} is used for pure model generalizability evaluation. Specifically, CIFAR-10 has 15k data points in the candidate pool of $\mathcal{D}_{\rm priv}$}, Fashion-MNIST and SVHN have 25k, and PathMNIST has 10,004. As for $\mathcal{D}_{\rm val}$}, its size is 10k for both CIFAR-10 and Fashion-MNIST, 26,032 for SVHN, and 7,180 for PathMNIST. $\mathcal{D}_{\rm infer}$} in Section IV-B has a size of $|{\mathcal{D}_{\rm protected}}|\cdot(|{\mathcal{D}_{\rm protected}}|-1)$ (around 2.2M). In Section IV-D, $|{\mathcal{D}_{\rm priv}}|$ is set to 125, 250, 500, 1k and 1.5k, whereas $|{\mathcal{D}_{\rm infer}}|$ is set to 15,500, 62,250, 250k and 500k, respectively. In this section, another set of comparison experiments is done with $|{\mathcal{D}_{\rm priv}}|$ fixed at 1.5k and $\mathcal{D}_{\rm infer}$} has a varying size of 250k, 500k, 1M, and 2M (i.e., all possible superimposition pairs, the same as Section IV-B). $\mathcal{M}_{\rm L}$}’s learning rate is 0.1 for Fashion-MNIST and 0.001 for SVHN and PathMNIST. The number of training epochs is set to 15 across all the datasets in all scenarios. For the latent space analysis, we train the VAE for 200 epochs with a learning rate set to 0.001 for Fashion-MNIST. The number of epochs is set to 300, and the learning rate is 0.0001 for SVHN and PathMNIST.
TABLE VIII: Sensitivity Analysis on $|{\mathcal{D}_{\rm priv}}|$ and $|{\mathcal{D}_{\rm infer}}|$ with ResNet-18 ($\mathcal{M}_{\rm L}$}) (Continued from Table III).